you have 3 $1, 1 $5 and 1 $10. draw 1 at a time without replacement until a $10 bill is drawn. What is the probability of winning $16. please show me how to work this out
Answer by Edwin McCravy(20067)



P{[($1 first) AND ($5 second) AND ($10 third)] OR [($5 first) and ($1 second) and ($10 third)]} =

(3/5)*(1/4)*(1/3) + (1/5)*(3/4)*(1/3) = 1/20 + 1/20 = 2/20 = 1/10

Edwin